OpenStudy (anonymous):

how far will a freely falling object fall from rest in five seconds ? six seconds ?

OpenStudy (anonymous):

any time there is a 'free-fall' question, it should be a trigger in your mind to use the 'kinematic equations' there are three of them like this: V = V + a*t V^2 = V^2 + 2*a*(s2-s1) S = S + vt+ (1/2)*a*t^2 the equations can casually be said like this: event 1 = event 2 meaning event 1 happens, then event 2 happens the side of the equation that has one term (the left side with either V^2, or V, or S) is the side which is 'event 2', the other side of the equation with all the terms is 'event 1' make sense so far?

OpenStudy (anonymous):

yup thanks !

OpenStudy (anonymous):

so, in the problem, we are told 3 things: 'free-fall' meaning a= gravity 'from rest' meaning V_initial = 0 and '5 seconds' meaning t = 5 now, knowing a, v and t, can you pick out which equation we could use to solve for how far an object will fall (how far an object will fall is the variable 'S')?

OpenStudy (anonymous):

I think its V= V+a*t

OpenStudy (anonymous):

close, but this equation would tell us the speed or velocity (V) at 5 seconds, we want to find how far it's fallen, so we want to find distance, so we will use the equation S = S + vt+ (1/2)*a*t^2 the left S is the amount fallen the right S is the starting height the v is for initial veloicty the t is for seconds the a is for acceleration (gravity)

OpenStudy (anonymous):

So I can help you simplify this S = S + vt+ (1/2)*a*t^2 S = how far it fell (answer we are looking for) S on the right side = 0, because we arn't on a platform or anything V = 0, becuase the object is initially at rest a = gravity, which is either 32.2 or 9.81 depending on if you are using feet or meters t = 5 seconds, becuase that is the time we are given